Given the supply function and inverse demand function, Qs = 40 + 20 P and p=10-1/10Q

respectively. Determine the equilibrium price and quantity?


1
Expert's answer
2021-04-29T10:28:14-0400

If p = 10 - 1/10Q, then Qd = 100 - 10p.

In equilibrium Qd = Qs, so:

100 - 10P = 40 + 20P,

30P = 60,

P = 2,

Q = 100 - 10×2 = 80 units.